OpenStudy (anonymous):

So I have to solve this stumper: triangle-square=circle square+triangle=2 circle circle*square=circle circle= square= triangle= It's due tomorrow!!!Please help.

OpenStudy (kkutie7):

If I were you I'd solve two equations for one thing... like circle for example and then plug the others into the one you didn't mess with... if that makes sense

OpenStudy (dangerousjesse):

Okay if circle\(\times\)square=circle then square=1

OpenStudy (dangerousjesse):

Or square could just = 0

OpenStudy (dangerousjesse):

Then all of them could just be 0.

OpenStudy (anonymous):

no they all can't be zero because the second sum is two digit

OpenStudy (dangerousjesse):

Or triangle=3 square=1 circle=2

OpenStudy (dangerousjesse):

It's not two digit.

OpenStudy (dangerousjesse):

It's 2\(\times\) circle.

OpenStudy (anonymous):

that's what the problem said

OpenStudy (kkutie7):

triangle-square=circle triangle-square=1 square+triangle=2 circle square+triangle=2(1) square+triangle=2 circle*square=circle square=circle/circle square=1

OpenStudy (anonymous):

but I didn't know who to type that so yeah

OpenStudy (dangerousjesse):

Can you copy and paste the entire thing?

OpenStudy (anonymous):

hold on

OpenStudy (triciaal):

I got the same result as @DangerousJesse

OpenStudy (anonymous):